using the head to tail method of vector addition, sketch the vector c given by c=3A-0.5B where A= 15lb due north B=40lb @35 degrees south of east.... Determine the magnitude and direction of C using the law of sines and the law of cosines
Answer #1

A = 15 j

B = 40 ( cos 35 i - sin 35 j) = 32.77 i - 22.943 j

vector C is given by

C = 3A - 0.5 B

C = 45 j - 16.385 i + 11.47 j

C = - 16.385 i + 56.47 j

magnitude

C = sqrt ( 16.385^2 + 56.47^2) = 58.8 lb

directio

x = arctan ( 56.47 / 16.388)

x = 73.82 North of west

========

Comment before rate in case any doubt, will reply for sure.. goodluck